FIND EACH PERCENT OF CHANGE. ROUND TO THE NEAREST TENTH. STATE WHETHER THE CHANGE IS AN INCREASE OR DECREASE.
1.) 144 to 204
2.) 758 to 604
3.) 8.3 to 9.10
I need this ASAP!
Thanks whoever solves this!

Answer by edjones(8007)

1.) 144 to 204
2.) 758 to 604
3.) 8.3 to 9.10
.
1) 204*100/144=144.7% increase.
2) 604*100/758=79.7% decrease
3) 9.1*100/8.3=109.6% increase
.
Ed